University of Niš, Faculty of Science and Mathematics

5th International Symposium on

MATHEMATICAL ANALYSIS AND ITS APPLICATIONS

MAA 5

NIŠKA BANJA, October, 2-6, 2002

First announcement

Following the tradition of the previous four symposia on Mathematical Analysis and its Applications, under the auspices of the Union of the Mathematical Societies of Yugoslavia, the Faculty of Science and Mathematics, University of Niš, organizes the Fifth Symposium in the spa of Niška Banja, 10 km from Niš, from Wednesday 2nd October to Sunday 6th October, 2002. The scientific program consists of lectures (45 minutes) and research talks (20 minutes). The program of the Symposium and the book of abstracts will be available at the Symposium. Submitted papers will be published in the Proceedings of the Symposium after the usual refereeing process.

Sections:

1. Functional Analysis and Operator Theory

2. Real and Complex Analysis

3. Differential Equations

4. Applications of Analysis

Pogram committee

V. Rakočević, Niš, Chair Z. Kadelburg, Beograd Ž. Pavićević, Podgorica S. Pilipović, Novi Sad

Organizing committee

D. Djordjević, Chair D. Cvetković S. Janković I. Jovanović M. Jovanović E. Malkowsky J. Manojlović M. Milosavljević S. Tričković S. Vujović

Conference information and contacts

MAA5 Organizing Committee Faculty of Science and Mathematics University of Niš Višegradska 33 18 000 Niš, Serbia, Yugoslavia

Fax: (+381 18) 533014 Phone: (+381 18) 324 492 E-mail: maa5@pmf.pmf.ni.ac.yu www.pmf.ni.ac.yu Important dates:

Deadline for the submission of abstracts: May 15, 2002 Preliminary registration: July 15, 2002 MAA 5 October 2-6, 2002

Registration form MAA 5

Name/Names

Institution

Address

E-mail

Phone

Fax

Section: 1 2 3 4

Hotel accommodation: Yes / No

The title of the talk

Abstract